OpenStudy (anonymous):

Find the exact value of cos 11π/3

OpenStudy (campbell_st):

well you are looking at a value which is pi/3 short of 4pi so its 4th quadrant, where cos is positive so find \[\cos(\frac{\pi}{3})\] you may need an exact ratio triangle.

OpenStudy (campbell_st):